According to Newton’s First law, accelerarion is said to be zero if and only if the force is zero...then..
  1. give an example where the vector sum of all forces is zero yet the particle is accelerated
  2. give an example where the vector sum of all forces is not zero yet the particle has acceleration

Grade:12th pass

1 Answers

Black panther
5 years ago
Let F be magnitude of vector sum of all force 
F=m (v-u)/t  
Now net force is 0 
So,
0=m (v-u)/t  
m (v-u)=0 
v-u=0 
v=u
So, accearetion=v-u/t
                           =v-v/t
                           =o
Hence your first condition is impossible 
Second case is general .
Any give body having vector sum of force non zero have accearetion 
Like hitting a cricket ball.